# MemeBlip2: A Novel Light Weight Multimodal System to Detect Harmful Memes thanks: :

### v1

@@ 5 Conclusion And Further Work
In this paper, we proposed MemeBLIP2, a lightweight yet effective multimodal system for the classification of harmful memes. By building upon the foundation established by MemeCLIP and integrating the BLIP-2 vision-language model, we introduced a modular architecture composed of modality-specific linear projection layers and adapter modules. These components align image and text representations into a shared semantic space while preserving computational efficiency. Evaluations on the PrideMM dataset demonstrate that MemeBLIP2 achieves competitive performance and excels in handling culturally nuanced and subtly ironic meme content. Our ablation studies underscore the critical role of the projection and adapter layers. The projection modules serve as semantic bridges between heterogeneous image and text embeddings, while the adapters refine feature representations through residual learning. The MLP-based classifier, equipped with nonlinear transformations, further enhances classification performance, particularly in scenarios involving class imbalance and complex semantic interactions. Unlike prior approaches that often rely on large-scale, resource-intensive architectures, MemeBLIP2 emphasizes architectural simplicity and efficiency. This design choice aligns with recent trends in developing compact multimodal models suitable for deployment in real-world scenarios, including environments with limited computational resources.
To enhance the deployability of MemeBLIP2 in real-time systems, future work will focus on model optimization techniques such as pruning, quantization, and knowledge distillation. These efforts aim to reduce inference latency and memory consumption, making the framework suitable for mobile and edge computing applications, including on-device content moderation. We also plan to broaden the applicability of MemeBLIP2 to multilingual and culturally diverse meme content. This includes integrating multilingual language models and expanding evaluation to cross-cultural datasets. Additionally, we intend to incorporate explainability mechanisms—such as attention visualization and gradient-based attribution—to improve the interpretability of the model’s decisions, which is crucial for transparency and ethical AI usage in content moderation.
@@
Acknowledgements
The authors conducted this research independently. The views and conclusions expressed in this work are solely those of the authors and do not necessarily reflect the official policy or position of Walmart.

### v3

@@ 5 Conclusion And Further Work
In this paper, we proposed MemeBLIP2, a lightweight but effective multimodal system for the classification of harmful memes. By building upon MemeCLIP and integrating the BLIP-2 vision-language model, we introduced a modular architecture composed of modality-specific linear projection layers and adapter modules. These components align image and text representations into a shared semantic space while preserving computational efficiency. Evaluations on the PrideMM dataset demonstrate that MemeBLIP2 achieves competitive performance and excels in handling culturally nuanced and subtly ironic meme content. Our ablation studies underscore the critical role of the projection and adapter layers. The projection modules serve as semantic bridges between heterogeneous image and text embeddings, while the adapters refine feature representations through residual learning. The MLP-based classifier, equipped with nonlinear transformations, further enhances classification performance, particularly in scenarios involving class imbalance and complex semantic interactions. Unlike prior approaches that often rely on large-scale, resource-intensive architectures, MemeBLIP2 emphasizes architectural simplicity and efficiency. This design choice aligns with recent trends in developing compact multimodal models suitable for deployment in real-world scenarios, including environments with limited computational resources. To enhance the deployability of MemeBLIP2 in real-time systems, future work will focus on model optimization techniques such as pruning, quantization, and knowledge distillation. These efforts aim to reduce inference latency and memory consumption, making the framework suitable for mobile and edge computing applications, including on-device content moderation. We also plan to broaden the applicability of MemeBLIP2 to multilingual and culturally diverse meme content. This includes integrating multilingual language models and expanding evaluation to cross-cultural datasets. Additionally, we intend to incorporate explainability mechanisms—such as attention visualization and gradient-based attribution—to improve the interpretability of the model’s decisions, which is crucial for transparency and ethical AI usage in content moderation.